OpenStudy (anonymous):

How do I expand 4cos(6x) in terms of sines using the double angle identity?

OpenStudy (anonymous):

Your attempt?

OpenStudy (anonymous):

Hint 6x = 2(3x)

OpenStudy (anonymous):

\(cos 2(x) = 1- 2sin^2 (x)\) now you have \(cos 2(3x) = 1-2sin^2(3x)\) , got me so far??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so?? just times4 in the front both sides, you get what you want to get. \(\color {red}{4}cos (6x)= \color{red}{4}*....\) can you finish??

OpenStudy (anonymous):

Yeah. Totally. That was way simpler than I was trying to make it be.

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

hahahahaa

OpenStudy (anonymous):

I was breaking it into cos(3x+3x) and using that identity and it was a total mess. Thank you.

OpenStudy (anonymous):

ok